1) FIND THE SOLUTION OF THE SYSTEM OF EQUATIONS: A COLLEGE BASKETBALL TEAM MADE 40 FIELD GOALS IN A RECENT GAME, SOME 2 POINTERS AND THE REST 3 POINTERS. ALL TOGETHER THE 40 BASKETS COUNTED FOR 89 POINTS. HOW MANY OF EACH TYPE OF FIELD GOALS WAS MADE?

2)SUPPOSE THAT 6.5 MILLION POUNDS OF COFFEE ARE SOLD WHEN THE PRICE IS $8 PER POUND AND 4.O MILLION POUNDS ARE SOLD WHEN IT IS $9 PER POUND.
(A) FIND A LINEAR FUNCTION THAT EXPRESSES THE AMOUNT OF COFFEE SOLD AS A FUNCTION OF THE PRICE PER POUND.
(B) USE THE FUNCTION OF PART A TO PEDICT THE NUMBER OF POUNDS OF COFFEE SOLD WHEN THE PRICE IS $6 PER POUND.

Answer by checkley71(8403)

1) 2X+3Y=89 & X+Y=40 THEN X=40-Y & 2(40-Y)+3Y=89 OR 80-2Y+3Y=89 OR
Y=9 THEN X=40-9 OR 31
PROOF 2*31+3*9=89 OR 62+27=89 AND 31+9=40
2) (6.5,8)(4,9) PROVIDES A SLOPE OF (9-8)/(4-6.5) OR 1/-1.5 WHICH SAYS THAT FOR EVERY DOLLAR INCREASE IN A POUND OF COFFEE YOU LOSE 1.5 MILLION POUNDS. THUS A DROP IN PRICE FROM 8 TO 6 POUNDS SHOULD INCREASE 2*1.5 MILLION SALES OR 6.5+3=9.5 MILLION POUNDS @ $6 PER POUND.